A reheat cycle with one stage of reheating is executed with steam initially from 90 bar and

530 C. The reheater pressure is 10 bar and the steam leaves the reheater at 480 C.

Condensation occurs at 0.03 bar. find a) W and Wp, b) Heat added, c) Heat rate of the engine


1
Expert's answer
2021-10-18T03:44:16-0400

Heat input in boiler


Qin = m (hi-hig) + in (hg-h₂) + (m-mpm₂) (his-hy) = 3.02 (3360-3-865:22) + 3-02 (3534-4-29317)


+ (3-01-0-532-0-0716) (2584-3-2167.2) = 10824-5 ks/ (kW).


Ideal Rankine ile Estruny Net work Cisentropie)


530+4-113-34


10824-5 20.479 (47.9%)


Actual Rankine Cycle efficiency actional mode


20.4448


4815-18


(44-48%)


& Plant Capacity - Actual Work (Power)


24815-1880-950-9 = 4116.99 kw


= 4.12 MW


(5) Heat released in condenser = = (3.02- 0.094 - 0.843) (2561.7- 251.12) 5442.2 KJ/s


Heat rejected by steam = Heat absorbed by boiling water = = mass × shc × ∆T


5442.2= mw × 4.187 × (21-10)


mass = 118.16 kg/s